# Exact Solutions to Acoustoelectric Interactions in Arbitrary Geometries

### v1

@@ Abstract
Acoustoelectric interactions occur when free carriers in a semiconductor interact with the fields of an acoustic wave in a piezoelectric medium. These interactions can amplify acoustic waves, as well as give rise to extremely large phononic nonlinearities and strong non-reciprocal effects. Despite the tremendous progress in the last ten years, the field is entirely dependent on analytical and perturbative solutions for the two simplest arrangements of piezoelectric-semiconductor materials. While these models have allowed the field to advance substantially, new geometries are arising that do not satisfy assumptions integral to these canonical models. These models rely on simplifying assumptions that remove the tensorial nature of the materials, restricting analysis to plane wave and perturbative solutions. Such restrictions fails to capture the non-perturbative nature of the acoustoelectric effect, illustrating the need for more advanced computational methods to analyze acoustoelectric systems. We develop, for the first time, a finite element method (FEM) model to solve for acoustoelectric interactions in arbitrary geometries. We use the model to verify existing results for amplification, dispersion, and non-reciprocity obtained from the canonical models. We then examine the acoustoelectric effect in two geometries not covered by the canonical models: a thin piezoelectric film placed above a semiconductor substrate and a fully 2D waveguide under a thin semiconductor layer. This work lays the foundation for accurate modeling of arbitrary acoustoelectric geometries such as those currently being developed for all-acoustic radio frequency (RF) signal processing, acoustoelectrically enhanced photonic devices, and quantum acoustoelectric devices.
@@ I Introduction
The acoustoelectric effect arises from the interactions between the fields of an acoustic wave in a piezoelectric medium and the free carriers inside of a semiconductor [28]. These interactions dramatically change the acoustic properties of a material resulting in large shifts in wave velocity [3, 27], attenuation and gain [9, 19], non-reciprocity [10, 20, 21], giant effective phononic nonlinearities [4, 8], and acoustoelectrically enhanced optomechanical interactions [24]. Approximate but accurate expressions for these effects have been derived in the case of bulk piezoelectric semiconductors, where the material platform is homogeneous [13, 4]. Additionally, a perturbation-based model has been derived for the separated medium amplifier, where the moving charge is separated from the piezoelectric material by a small air gap and coupled to the acoustic wave via the electric field [18, 16]. The geometries for these two canonical models are shown in Figure 1. Despite the fact that these canonical models have allowed accurate predictions of acoustoelectric devices for amplification [17], radiofrequency (RF) switching [25], nonlinear mixing [22], and phononic oscillators [26], they are difficult to extend beyond these simple geometries. The homogeneous model, while admitting an exact solution, ignores the tensorial nature of the material so as to focus on a single polarization of plane waves; meanwhile, the perturbation-based model makes many approximations and cannot capture any elastic coupling that occurs when the semiconductor is in direct contact with the piezoelectric substrate, which is the most common scenario in real state-of-the-art devices. This highlights the limitations on calculating the exact acoustoelectric effects for real-world systems, for which the canonical models fail to maintain their accuracy.
@@
More comprehensive and arbitrary models are needed to accurately predict device performance in novel geometries including, but not limited to, thin piezoelectric substrates placed on top of a semiconductor layer, spatially dependent carrier concentration in the semiconductor, phononic waveguides with strong horizontal and vertical confinement, and phononic cavities, a subset of which are illustrated in Figure 2. Extending acoustoelectric devices to these geometries is vital to optimizing acoustoelectric interactions, understanding electrical boundary conditions, controlling dispersion, and refining device architecture. It would be expected that the field of acoustoelectric devices would progress much faster if powerful and automatable tools, such as finite element method (FEM) modeling, were amenable to device design. As such, we develop and evaluate an FEM model that incorporates acoustoelectric interactions to solve for the acoustic modes and their corresponding voltage-dependent gain and dispersion curves in the presence of a DC-bias drift field.
@@ II The Acoustoelectric Interaction
Figure 2: a) Platform with piezoelectric material on top of a substrate that has been doped to have carriers available for the acoustoelectric effect. This structure is not covered by the canonical models due elastic coupling between the semiconductor and piezoelectric layers. b) Piezoelectric waveguides on a substrate with the semiconductor on top. The canonical models do not take into account rapid transverse horizontal variations in field strength introduced by the acoustic waveguide. c) Optomechanical cavity utilizing the acoustoelectric effect to alter acoustic breathing modes to control phononic properties for enhancement of Brillouin scattering. Such an arbitrary geometry is not covered by existing models.
@@ IV Rayleigh Waves in the Separated Medium Amplifier
We compare Equations 22 and 23 for a finite-thickness semiconductor to our FEM model in Figure 4. $Y^{-}$ for a finite-thickness semiconductor has been previously derived and is included in the supplementary information for completeness, as well as $Y^{-}$ for the infinitesimally thin and semi-infinitely thick cases [16]. The structure of the separated medium model is shown in Figure 4a, where we utilize $Y$ cut, $Z$ propagating lithium niobate (LiNbO3). The mode displacement, $|\mathbf{u}|$, and the longitudinal electric field, $E_{x}$, for the Rayleigh mode are shown in Figures 4b and 4c, respectively. It is worth pointing out that we can observe screening effects on the electric field as evidenced by the phase shift of the field above the piezoelectric layer induced by the free carriers within the semiconductor layer as seen in the inset image of Figure 4c which is taken edge on with the surface parallel to the $xz$-plane.
Sweeping the drift field for a Rayleigh mode at 1 GHz, our model is in close agreement to the gain curve obtained from Equation 22, as shown in Figure 4d. In Figure 4e, for the same mode, we demonstrate the shift in the propagation constant is also in close agreement between our FEM model and Equation 23. For a fixed drift velocity of $v_{d}=3v_{a}$, we show the gain from our FEM model closely follows the results obtained from Equation 22 over a broad frequency range spanning from 20 MHz to 10 GHz in Figure 4f. The inset image shows the regime of comparison from the original work and includes experimental data points extracted from the original paper used to initially verify the model [16]. From this, we conclude that our FEM model is in agreement with the second canonical model for the case of a separated medium amplifier.
@@ VI Shear Horizontal Waves in a Fully 2D Waveguide
To understand how variation in the transverse horizontal direction of the field affects the calculated acoustoelectric gain, we turn to a waveguide geometry utilizing a ridge waveguide in Y-cut, X-propagating lithium niobate (LiNbO3) on silicon (Si) as shown in Figure 6a.The assymetry of the mode results form the strong anisotropy of the lithium niobate. We include a sidewall angle of 78 degrees take into account realistic fabrication constraints. A $50$ nm layer of indium gallium arsenide (In0.53Ga0.47As) is placed directly on top of the waveguide, providing access to carriers for the acoustic wave. We desire a fixed frequency of $f=4.7GHz$ when the acoustoelectric effect is turned off in the FEM model. This is to reduce the parameters to be varied when comparing to the perturbative approximation and to provide a better comparison between different waveguide widths. As such, we vary the wavelength as a function of waveguide width, $w$, to account for the shift of the dispersion relation. This results in a span of wavelengths ranging from $\lambda=728.8$ nm at $w=375$ nm to $\lambda=762$ nm at $w=750$ nm. For these parameters we find a shear horizontal mode as shown in Figures 6b and 6c.
@@ VII Conclusion
VIII Supplementary Information
VIII.1 Admittances

### v2

@@ Abstract
Acoustoelectric interactions occur when free carriers in a semiconductor interact with the fields of an acoustic wave in a piezoelectric medium. These interactions can amplify acoustic waves, as well as give rise to extremely large phononic nonlinearities and strong non-reciprocal effects. The field of acoustoelectric devices is currently dependent on analytical and perturbative solutions for the two simplest arrangements of piezoelectric-semiconductor materials. While these canonical models have allowed the field to advance substantially, new geometries are arising that do not satisfy assumptions integral to these models. These assumptions include the treatment of the interactions between the acoustic fields and free carriers as weak, the neglect of the tensorial nature of the material properties, the omission of the spatial variations in the phonons’ electric field profiles, and the disregard of elastic coupling across material boundaries, among others. We develop, for the first time, a finite element method (FEM) model to solve for acoustoelectric interactions in arbitrary geometries that avoids making the assumptions of the canonical models. We verify the FEM model using results for amplification, dispersion, and non-reciprocity obtained from the canonical models in their regime of validity. We then examine the acoustoelectric effect in two geometries not covered by the canonical models: a thin piezoelectric film placed on a semiconductor substrate and a fully 2D waveguide under a thin semiconductor layer. This work lays the foundation for accurate modeling of arbitrary acoustoelectric geometries such as those currently being developed for all-acoustic radio frequency (RF) signal processing, acoustoelectrically enhanced photonic devices, and quantum acoustoelectric devices.
@@ I Introduction
The acoustoelectric effect arises from the interactions between the fields of an acoustic wave in a piezoelectric medium and the free carriers inside of a semiconductor [29]. These interactions dramatically change the acoustic properties of a material resulting in large shifts in wave velocity [3, 28], attenuation and gain [9, 20], non-reciprocity [10, 21, 22], giant effective phononic nonlinearities [4, 8], and acoustoelectrically enhanced optomechanical interactions [25]. Approximate but accurate expressions for these effects have been derived in the case of bulk piezoelectric semiconductors, where the material platform is homogeneous [13, 4]. Additionally, a perturbation-based model has been derived for the separated medium amplifier, where the moving charge is separated from the piezoelectric material by a small air gap and coupled to the acoustic wave via the electric field [18, 16]. The geometries for these two canonical models are shown in Figures 1a and 1b, respectively. Despite the fact that these canonical models have allowed accurate predictions of acoustoelectric devices for amplification [17], radiofrequency (RF) switching [26], nonlinear mixing [23], and phononic oscillators [27], they are difficult to extend beyond these simple geometries. The homogeneous model, while admitting an exact solution, ignores the tensorial nature of the material so as to focus on a single polarization of plane waves; meanwhile, the perturbation-based model makes many approximations and cannot capture any elastic coupling that occurs when the semiconductor is in direct contact with the piezoelectric substrate, which is the most common scenario in real state-of-the-art devices. This highlights the limitations on calculating the exact acoustoelectric effects for real-world systems, for which the canonical models fail to maintain their accuracy.
@@
More comprehensive and arbitrary models are needed to accurately predict device performance in novel geometries including, but not limited to, thin piezoelectric films placed on top of a semiconductor substrate, spatially dependent carrier concentration in the semiconductor, phononic waveguides with strong horizontal and vertical confinement, and phononic cavities, a subset of which are illustrated in Figure 2. In particular, complicated geometries, such as the acoustoelectrically enhanced optomechanical cavity shown in Figure 2c, have no analog in the canonical models [19]. Extending acoustoelectric models to these geometries is vital to optimizing acoustoelectric interactions, understanding electrical boundary conditions, controlling dispersion, and refining device architecture. It would be expected that the field of acoustoelectric devices would progress much faster if powerful and automatable tools, such as finite element method (FEM) modeling, were amenable to device design. As such, we develop and evaluate an FEM model that incorporates acoustoelectric interactions to solve for the acoustic modes and their corresponding voltage-dependent gain and dispersion curves in the presence of a DC-bias drift field.
@@ II The Acoustoelectric Interaction
Figure 2: a) Platform with piezoelectric material on top of a substrate that has been doped to have carriers available for the acoustoelectric effect. This structure is not covered by the canonical models due elastic coupling between the semiconductor and piezoelectric layers. b) Piezoelectric waveguides on a substrate with the semiconductor on top. The canonical models do not take into account rapid transverse horizontal variations in field strength introduced by the acoustic waveguide. c) Optomechanical cavity utilizing the acoustoelectric effect to alter acoustic breathing modes to control phononic properties for enhancement of Brillouin scattering [19]. Such an arbitrary geometry is not covered by existing models.
@@ IV Rayleigh Waves in the Separated Medium Amplifier
We compare Equations 22 and 23 for a finite-thickness semiconductor to our FEM model in Figure 4. $Y^{-}$ for a finite-thickness semiconductor has been previously derived and is included in the appendix for completeness, as well as $Y^{-}$ for the infinitesimally thin and semi-infinitely thick cases [16]. The structure of the separated medium model is shown in Figure 4a, where we utilize $Y$ cut, $Z$ propagating lithium niobate (LiNbO3). The mode displacement, $|\mathbf{u}|$, and the longitudinal electric field, $E_{x}$, for the Rayleigh mode are shown in Figures 4b and 4c, respectively. It is worth pointing out that we can observe screening effects on the electric field as evidenced by the phase shift of the field above the piezoelectric layer induced by the free carriers within the semiconductor layer as seen in the inset image of Figure 4c which is taken edge on with the surface parallel to the $xz$-plane.
Sweeping the drift field for a Rayleigh mode at 1 GHz, our model is in close agreement to the gain curve obtained from Equation 22, as shown in Figure 4e. In Figure 4f, for the same mode, we demonstrate the shift in the propagation constant is also in close agreement between our FEM model and Equation 23. For a fixed drift velocity of $v_{d}=3v_{a}$, we show the gain from our FEM model closely follows the results obtained from Equation 22 over a broad frequency range spanning from 20 MHz to 10 GHz in Figure 4d. The inset image shows the regime of comparison from the original work and includes experimental data points extracted from the original paper used to initially verify the model [16]. From this, we conclude that our FEM model is in agreement with the second canonical model for the case of a separated medium amplifier.
@@ VI Shear Horizontal Waves in a Fully 2D Waveguide
To understand how variation in the transverse horizontal direction of the field affects the calculated acoustoelectric gain, we turn to a waveguide geometry utilizing a ridge waveguide in Y-cut, X-propagating lithium niobate (LiNbO3) on silicon (Si) as shown in Figure 6a. The asymmetry of the mode results form the strong anisotropy of the lithium niobate. We include a sidewall angle of 78 degrees take into account realistic fabrication constraints. A $50$ nm layer of indium gallium arsenide (In0.53Ga0.47As) is placed directly on top of the waveguide, providing access to carriers for the acoustic wave. We desire a fixed frequency of $f=4.7GHz$ when the acoustoelectric effect is turned off in the FEM model. This is to reduce the parameters to be varied when comparing to the perturbative approximation and to provide a better comparison between different waveguide widths. As such, we vary the wavelength as a function of waveguide width, $w$, to account for the shift of the dispersion relation. This results in a span of wavelengths ranging from $\lambda=728.8$ nm at $w=375$ nm to $\lambda=762$ nm at $w=750$ nm. For these parameters we find a shear horizontal mode as shown in Figures 6b and 6c.
@@ VII Conclusion
Appendix A Admittances